OpenStudy (chaotic_butterflies):

Trig help? How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? - more information below

OpenStudy (phi):

I would delete this part from your answer to Part 1 ***Using the information presented earlier, just multiply the 6 degrees by 25 minutes. 6*25 = 150, so the minute hand moves 150 radians from 1:25 to 1:50*** what you have below that line is correct: you get 6*25= 150 degrees and that is 5 pi/6 radians

OpenStudy (phi):

For Part 2, this is muddled ***2pi(4)(5pi/6)/2pi = 10pi /3 inches**** you want arc= radius * angle(in radians) so just 4 * 5pi/6 = 10pi/3 inches (about 10.5 inches)

OpenStudy (chaotic_butterflies):

Oh, okay @Phi

OpenStudy (phi):

For the posted question **How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? ** use radius * angle = arc length they tell you radius is 4 they tell you arc length is 5 pi put those numbers into the formula what do you get ?

OpenStudy (chaotic_butterflies):

4*a = 5pi a = 5pi/4

OpenStudy (phi):

yes, so the angle is 5 pi/4 radians (or 225 degrees) It will be halfway between 37 and 38 minutes past the hour

OpenStudy (chaotic_butterflies):

Awesome... so how do I find the coordinate point associated with that radian measure? I haven't even heard of finding something like that before .~.

OpenStudy (phi):

exactly what is the question they asked?

OpenStudy (chaotic_butterflies):

Here's the whole thing: Part 1: How many radians does the minute hand move from 1:25 to 1:50? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (phi):

I guess we put the origin at the center of the clock